OpenStudy (twizttiez):

What is the volume of a sphere that has a diameter of 18 in.? Use 3.14 to approximate pi and express your answer in hundredths. in3

OpenStudy (wolf1728):

Sphee Volume = 4/3 *PI*r^3

OpenStudy (hockeychick23):

you can use the same formula i gave you for the last one, you just divide the diameter by 2 to get the radius which is what we used for the last problem: 4/3πr^3 V=3053.63

OpenStudy (wolf1728):

Volume = 4/3 * 3.14* 9 * 9 * 9 Volume = 3,052.08

OpenStudy (wolf1728):

I think hockeychick used the full value of PI 3.14159265358... whereas I used 3.14

OpenStudy (twizttiez):

Well thank you both

OpenStudy (hockeychick23):

anytime :)

OpenStudy (wolf1728):

You were supposed to use 3.14

OpenStudy (wolf1728):

u r welcome

OpenStudy (wolf1728):

Thanks for the medal Twitz